# thanks: Corresponding author

### v1

@@ III.1 Inequality for the probabilities of projective measurements in MEBs of a qudit system
$f_{L}(q)=1+(L-1)\max_{k}|N_{k,d-k}|.$ (43)
Here, $|N_{k,d-k}|$ is given by
where $\mu=\sum_{j=0}^{d-1}q_{j}^{2}$ is defined as the index of coincidence of the distribution $q=(q_{0},\cdots,q_{d-1})$, and the maximum is taken over $k\in\{1,\cdots,(d-1)/2\}$ when $d$ is odd, and over $k\in\{1,\cdots,d/2\}$ when $d$ is even. Obviously, $|N_{k,d-k}|=0$ for MUBs for which $\mu=1/d$, thus leading to the results of MolmerPRA2009 .
@@
Building on the method presented in Ref. MolmerPRA2009 , we consider the composite system $\mathcal{H}\otimes\mathcal{H}$ and define the following normalized, but not necessarily orthogonal, vectors
@@
where $\omega={\mathrm{e}}^{2\pi i/d}$, $k=1,\cdots,d-1$ and $\alpha=1,\cdots,L$. Here, $|e_{j}^{(\alpha)}\rangle^{\ast}$ denotes the complex conjugate of $|e_{j}^{(\alpha)}\rangle$ with respect to a definite basis, say the standard basis $\{|e_{j}^{(1)}\rangle\}_{j=0}^{d-1}$. With this choice, the first basis is real, i.e., $|e_{j}^{(1)}\rangle^{\ast}=|e_{j}^{(1)}\rangle$ for $j=0,\cdots,d-1$, however, for all other bases the expansion coefficients differ by a complex conjugation when they are linear expanded in terms of the first basis. By definition, all these states are maximally entangled states of the two-qudit system $\mathcal{H}\otimes\mathcal{H}$. However, in general, they are neither orthogonal nor complete. More precisely, although the first vector is orthogonal to every vector of the second class, i.e., $\langle\phi|\phi_{k}^{(\alpha)}\rangle=0$, orthogonality fails to hold between vectors of the second class. This can be seen from the overlap or Gram matrix $G$ defined by
where $N_{kk^{\prime}}=\frac{1}{d}\sum_{j,j^{\prime}=0}^{d-1}\omega^{k^{\prime}j^{\prime}-kj}q_{j\oplus j^{\prime}}$ for $k,k^{\prime}=1,\cdots,d-1$ form the entries of the Gram submatrix $N$ for $\alpha\neq\alpha^{\prime}$, independent of $\alpha$ and $\alpha^{\prime}$. Noteworthy, the matrix $N$ is anti-diagonal as its nonzero entries can be expressed as
$N_{kk^{\prime}}=\begin{cases}\sum_{m=0}^{d-1}\omega^{k^{\prime}m}q_{m}&\mbox{if }k\oplus k^{\prime}=d\\ 0&\mbox{otherwise}.\end{cases}$ (48)
Obviously, $N_{kk^{\prime}}=0$ for $q_{k\oplus k^{\prime}}=1/d$, therefore the vectors defined in Eq. (46) remain mutually orthogonal when the underlying bases are MUBs.
Crucial to our analysis is the fact that the vectors given by Eqs. (45) and (46) are linearly independent, allowing them to span a $D$-dimensional subspace of $\mathcal{H}\otimes\mathcal{H}$ where D=L(d1)+1D=L(d-1)+1. Defining $\{|\chi_{l}\rangle\}_{l=1}^{d^{2}-D}$ as an orthonormal basis for the orthogonal complement of this subspace, and adding these vectors to the previously defined set of vectors we obtain, in general, a nonorthogonal basis for $\mathcal{H}\otimes\mathcal{H}$ as $\{|\phi\rangle,|\phi_{k}^{(\alpha)}\rangle,|\chi_{l}\rangle\}$ where $k=1,\cdots,d-1$ and $\alpha=1,\cdots,L$, and $l=1,\cdots,d^{2}-D$.
Now, for a given qudit quantum state $\varrho$, we define a two-qudit pure state $\varrho\otimes{\mathbb{I}}]|\phi\rangle$, and expand it in terms of the aforementioned basis of $\mathcal{H}\otimes\mathcal{H}$, we get
$[\varrho\otimes{\mathbb{I}}]|\phi\rangle=\frac{1}{d}|\phi\rangle+\sum_{\alpha=1}^{L}\sum_{k=1}^{d-1}a_{k}^{(\alpha)}|\phi_{k}^{(\alpha)}\rangle+|\chi\rangle,$ (49)
where the coefficients $a_{k}^{(\alpha)}$ are given by $a_{k}^{(\alpha)}=[G^{-1}\Omega]_{k}^{(\alpha)}$. A proof can be found in Appendix A. Here, $G^{-1}$ refers to the inverse of the Gram matrix $G$ defined by Eq. (III.1), and the L(d1)L(d-1)-dimensional vector $\Omega=[\Omega_{l}^{(\beta)}]$ is given by
@@
for $\beta=1,\cdots,L$ and $l=1,\cdots,d-1$. Based on this, the squared norm of the vector defined by Eq. (49) can be written as
@@
where the last line follows from the fact that the Gram matrix of a linearly independent set of vectors is positive definite, as such $\lambda_{\max}(G)=1/\lambda_{\min}(G^{-1})$. To complete the proof, it remains only to determine the eigenvalues of the Gram matrix $G$. A complete derivation of the eigenvalues is provided in Appendix B—from which the maximum eigenvalue we need here is given by $\lambda_{\max}(G)=f_{L}(q)$, where $f_{L}(q)$ is defined by Eq. (43). This concludes the proof. ∎
An important feature that follows from the proof is present here; the Gram matrix $G$, so the submatrix $N$, provides all information on the MEBs, meanwhile, the vector $\Omega\in\mathbb{R}^{d(L-1)}$ contains the probabilities of obtaining different outcomes when the system—initially prepared in state $\varrho$—is measured in MEBs. This feature enables us to parameterize the probability distribution $q=(q_{0},\cdots,q_{d-1})$ associated with each MEBs, using d1d-1 independent parameters provided by the entries of the matrix $N$. See Appendix C.
@@
where $K=\lfloor\frac{L}{C}\rfloor$, and $C$ is the upper bound for $\sum_{\alpha=1}^{{L}}\sum_{j=0}^{d-1}\big{[}p_{j}^{(\alpha)}\big{]}^{2}$. Here for MEBs, $C=f_{L}(q)\big{[}{\textrm{Tr}}[\varrho^{2}]-\frac{1}{d}\big{]}+\frac{L}{d}$ is the upper bound given in Eq. (42).
@@ IV Entanglement detection
where $q_{\max}=\max_{k}\{q_{0},\cdots,q_{d-1}\}$ and $f_{L}(q)$ is defined by Eq. (43). Also, $\mathcal{O}^{(b)}$ is a set of orthogonal rotation in $\mathbb{R}^{d}$ around the axis $\boldsymbol{n}_{\ast}=(1,1,\cdots,1)/\sqrt{d}$, so that $\mathcal{O}^{(b)}\boldsymbol{n}_{\ast}=\boldsymbol{n}_{\ast}$.
@@
$=$ $\frac{(d-1)}{d}L\gamma\omega,$
where in the the first equality we have used Eq. (2), and the second equality follows from the fact that the optimum detection happens when ${\textrm{Tr}}[\mathcal{O}^{(\alpha)}]$ takes its maximum value $d$, i.e., when $\mathcal{O}^{(\alpha)}={\mathbb{I}}_{d}$ for all $\alpha$. It turns out that $W$ can detect entanglement of the isotropic states when $\omega>\frac{1}{L\gamma}$. Figure 5 illustrates the behaviour of this function in terms of the index $\mu$ for $d=3$ and different values of $L=2,3,4$. In all cases, the solid curves denote the bounds when MEBs are used to construct witnesses. The dashed lines, on the other hand, captures the bounds when a set of $L$ MUBs are used to construct witnesses. Not surprisingly, for a fixed $L$ the MUBs-based witness is finer than the MEBs-based witness, however, if a MEBs-based witness employs a larger set of measurements, it can surpass the MUBs-based witness in detecting entanglement, particularly within the specific index $\mu$ of the underlying distribution $q$ of the MEBs.
Figure 5: Plots of $(L\gamma)^{-1}$ in terms of index $\mu$ for $d=3$ and different values of $L=2,3,4$. For equal $L$s, the MUBs-based witness (dashed lines) always captures a wider range of states relative to the MEBs-based witness (solid curves). However, if a MEBs-based witness employs a larger set of measurements, it can surpass the MUB-based witness in detecting entanglement, particularly within low values of the index $\mu$.
@@ V Conclusion
In this work, we introduced the concept of mutually equi-biased bases (MEBs), a generalization of the mutually unbiased bases (MUBs). For such bases, we relaxed the condition of mutual maximum unpredictability that holds for MUBs; instead, we associated a probability distribution $q=(q_{0},\cdots,q_{d-1})$ with each of the MEBs. This distribution captures the predictability of the outcomes of a measurement in one basis when the system is prepared in a state from the other basis. However, the information extracted from the measurement is unbiased, meaning that it cannot reveal which state from which basis the system is prepared in, and which basis is used for the measurement.
We derived a complete set of MEBs for $d=2$ and $d=3$, and found that although for $d=2$ a complete set of three MEBs exist for all values of the index $\mu=\sum_{k=0}^{d-1}q_{k}^{2}$, for $d\geq 3$ this is not the case. More precisely, for $d=3$, the constraints imposed by the mutual equi-biasedness limits the range of the index $\mu$ to the feasible region $1/3\leq\mu\leq 1/2$ within which a set of four MEBs exists. Furthermore, in addition to this restriction on $\mu$, for a given $\mu$ in the feasible region the components of the probability distribution $q=(q_{0},q_{1},q_{2})$ are also limited to some specific discrete points; only one probability distribution qq—up to permutation of its components—can generate a complete set of four MEBs.
For a qudit system, we derived an inequality for the probabilities of projective measurements in MEBs. We further used the inequality to derive an entropic uncertainty inequality in MEBs. As expected, the maximal incompatibility inherent to MUBs breaks down in the case of MEBs of the same size. The inequality for probabilities is then used to obtain a class of positive maps, from which we constructed entanglement witnesses.
@@
These results provide a deeper understanding of the structure and properties of MEBs, offering new tools for entanglement detection and quantum measurement theory. The challenging dimension $d=6$ is currently under investigation to determine whether there exist values of the index $\mu$ for which one can construct a set of at least four MEBs—exceeding a set of three MUBs currently known for this dimension. We hope that our results will have broader applications in quantum information theory; in particular, for dimensions where no more than three MUBs are known, even a set of four MEBs with $\mu>1/d$ could shed light on the challenging open problem of constructing complete set of MUBs.

### v2

@@ III Entropic uncertainty relation for MEBs
[…]
where $\mu=\sum_{j=0}^{d-1}q_{j}^{2}$ is the index of coincidence of the distribution $q$. Define $\lambda_{\max}(N)=\max_{k}|N_{k,d-k}|$, where the maximum is taken over $k\in\{1,\ldots,(d-1)/2\}$ when $d$ is odd, and over $k\in\{1,\ldots,d/2\}$ when $d$ is even.
Let us now state the following lemma, which is essential for the remainder of this section.
Lemma 4.
For a two-qudit composite system $\mathcal{H}\otimes\mathcal{H}$, consider the following set of vectors
where $\omega={\mathrm{e}}^{2\pi i/d}$, $k=1,\ldots,d-1$, and $\alpha=1,\ldots,L$. Here, $|e_{j}^{(\alpha)}\rangle^{\ast}$ denotes the complex conjugate of $|e_{j}^{(\alpha)}\rangle$ with respect to a definite basis, say the standard basis $\{|e_{j}^{(1)}\rangle\}_{j=0}^{d-1}$. Then
  1. (a)
    The set of vectors (48) is linearly independent and span an L(d1)L(d-1)-dimensional subspace of $\mathcal{H}\otimes\mathcal{H}$ if and only if
    $\lambda_{\max}(N)|<1/(L-1).$ (49)
  2. (b)
    When (a) holds, the reciprocal vector $|\widetilde{\phi}_{k}^{(\alpha)}\rangle$, corresponding to each vector $|\phi_{k}^{(\alpha)}\rangle$, is defined by
    $|\widetilde{\phi}_{k}^{(\alpha)}\rangle=\sum_{k^{\prime}=1}^{d-1}\sum_{\alpha^{\prime}=1}^{L}[G^{-1}]_{k^{\prime}k}^{(\alpha^{\prime}\alpha)}|\phi_{k^{\prime}}^{(\alpha^{\prime})}\rangle,$ (50)
    where $G^{-1}$ is the inverse of the Gram matrix $G$, defined by $G_{kk^{\prime}}^{(\alpha\alpha^{\prime})}=\langle\phi_{k}^{(\alpha)}|\phi_{k^{\prime}}^{(\alpha^{\prime})}\rangle$, and reciprocity means $\langle\phi_{k}^{(\alpha)}|\widetilde{\phi}_{k^{\prime}}^{(\alpha^{\prime})}\rangle=\delta_{kk^{\prime}}\delta^{(\alpha\alpha^{\prime})}$.
Proof.
To prove part (a), consider the Gram matrix $G$, defined by
where $N_{kk^{\prime}}$ are entries of the matrix $N$, as
$N_{kk^{\prime}}=\frac{1}{d}\sum_{j,j^{\prime}=0}^{d-1}\omega^{k^{\prime}j^{\prime}-kj}q_{j\oplus j^{\prime}},$ (52)
for $k,k^{\prime}=1,\ldots,d-1$. Clearly, $N$ is anti-diagonal and its nonzero entries can be simplified as Eq. (4). In Appendix A, we derive the eigenvalues of $G$; in particular, its smallest and largest eigenvalues are $\lambda_{\min}(G)=1-(L-1)\lambda_{\max}(N)$ and $\lambda_{\max}(G)=1+(L-1)\lambda_{\max}(N)$, respectively. It follows therefore that the set of vectors (48) is linearly independent iff the Gram matrix is positive definite, i.e., $\lambda_{\min}(G)>0$, thereby completing the proof of part (a).
To prove part (b) of the lemma, suppose vectors (48) are linearly independent, i.e., $\lambda_{\max}(N)<1/(L-1)$. The reciprocal vectors can be expanded in terms of the original vectors as
$=$ $\left[CG^{\textrm{T}}\right]_{kl}^{(\alpha\beta)}.$
This implies that $CG^{\textrm{T}}={\mathbb{I}}_{L(d-1)}$, or equivalently, $C^{\textrm{T}}=G^{-1}$, where T stands for the transposition. Accordingly, up to a transposition, the matrix of the coefficients is obtained by the inverse of the Gram matrix $G$. This completes the proof of the lemma. ∎
@@ III.1 Inequality for the probabilities of projective measurements in MEBs of a qudit system
$f_{L}(q)=1+(L-1)\lambda_{\max}(N).$ (56)
Obviously, $|N_{k,d-k}|=0$ for MUBs for which $\mu=1/d$, thus leading to the results of MolmerPRA2009 .
@@
Building on the method presented in Ref. MolmerPRA2009 , we define a two-qudit pure state $\varrho\otimes{\mathbb{I}}]|\phi\rangle$, where $\varrho$ is an arbitrary quantum state on $\mathcal{H}$, and
@@
is a maximally entangled state of the composite system $\mathcal{H}\otimes\mathcal{H}$. One can easily show that $\langle\phi|\phi_{k}^{(\alpha)}\rangle=0$, for $k=1,\ldots,d-1$ and $\alpha=1,\ldots,L$. Invoking this and Lemma 4, vectors $\{|\phi\rangle,|\phi_{k}^{(\alpha)}\rangle\}$ span a $D$-dimensional subspace of $\mathcal{H}\otimes\mathcal{H}$ where D=L(d1)+1D=L(d-1)+1. Defining $\{|\chi_{l}\rangle\}_{l=1}^{d^{2}-D}$ as an orthonormal basis for the orthogonal complement of this subspace, and adding these vectors to the previously defined set of vectors we obtain, in general, a nonorthogonal basis for $\mathcal{H}\otimes\mathcal{H}$ as $\{|\phi\rangle,|\phi_{k}^{(\alpha)}\rangle,|\chi_{l}\rangle\}$ where $k=1,\ldots,d-1$ and $\alpha=1,\ldots,L$, and $l=1,\ldots,d^{2}-D$.
By expanding $\varrho\otimes{\mathbb{I}}]|\phi\rangle$ in terms of the aforementioned basis of $\mathcal{H}\otimes\mathcal{H}$, we get
$[\varrho\otimes{\mathbb{I}}]|\phi\rangle=\frac{1}{d}|\phi\rangle+\sum_{\alpha=1}^{L}\sum_{k=1}^{d-1}[G^{-1}\Omega]_{k}^{(\alpha)}|\phi_{k}^{(\alpha)}\rangle+|\chi\rangle,$ (58)
where the first term follows from $\langle\phi|[\varrho\otimes{\mathbb{I}}]|\phi\rangle=1/d$, and the unnormalized vector $|\chi\rangle$ represents the component of $\varrho\otimes{\mathbb{I}}]|\phi\rangle$ that lies in the orthogonal subspace spanned by $\{|\chi_{l}\rangle\}$. For the second term, using Lemma 4, we have $\langle\widetilde{\phi}_{k}^{(\alpha)}|[\varrho\otimes{\mathbb{I}}]|\phi\rangle=[G^{-1}\Omega]_{k}^{(\alpha)}$, where $G^{-1}$ refers to the inverse of the Gram matrix $G$ defined by Eq. (III). The L(d1)L(d-1)-dimensional vector $\Omega$ is defined by its components $\Omega_{l}^{(\beta)}$, as
@@
[…]
For the squared norm of the vector defined by Eq. (58), we have
@@
where the last line follows from the fact that the Gram matrix of a linearly independent set of vectors is positive definite, as such $\lambda_{\max}(G)=1/\lambda_{\min}(G^{-1})$. To complete the proof, it remains only to determine the eigenvalues of the Gram matrix $G$. A complete derivation of the eigenvalues is provided in Appendix A—from which the maximum eigenvalue we need here is given by $\lambda_{\max}(G)=f_{L}(q)$, where $f_{L}(q)$ is defined by Eq. (56). This concludes the proof. ∎
In proving the theorem, we observe two different features of the d(L1)d(L-1)-dimensional vector $\Omega$, and the $(d-1)\times(d-1)$ matrix $N$; The vector $\Omega$ contains the probabilities of obtaining different outcomes when the system—initially prepared in state $\varrho$—is measured in MEBs, meanwhile, the Gram matrix $G$, so the submatrix $N$, provides all information on the MEBs. This last feature enables us to parameterize the probability distribution $q=(q_{0},\ldots,q_{d-1})$ associated with each MEBs, using d1d-1 independent parameters provided by the entries of the matrix $N$, as $N_{k,d-k}=\delta_{2k-1}+i\delta_{2k}$, where (d1)(d-1) real parameters $\delta_{1},\ldots,\delta_{d-1}$ are defined by Eqs. (5) and (6), respectively, (see Sec. II.1.2).
@@
where $K=\lfloor\frac{L}{C}\rfloor$, and $C$ is the upper bound for $\sum_{\alpha=1}^{{L}}\sum_{j=0}^{d-1}\big[p_{j}^{(\alpha)}\big]^{2}$. Here, for a set of $L$ MEBs that satisfy the condition $\lambda_{\max}(N)<1/(L-1)$, $C=f_{L}(q)\big[{\textrm{Tr}}[\varrho^{2}]-\frac{1}{d}\big]+\frac{L}{d}$ is the upper bound given in Eq. (55).
@@ IV Entanglement detection
where $q_{\max}=\max_{k}\{q_{0},\ldots,q_{d-1}\}$ and $f_{L}(q)$ is defined by Eq. (56). Also, $\mathcal{O}^{(b)}$ is a set of orthogonal rotation in $\mathbb{R}^{d}$ around the axis $\boldsymbol{n}_{\ast}=(1,1,\ldots,1)/\sqrt{d}$, so that $\mathcal{O}^{(b)}\boldsymbol{n}_{\ast}=\boldsymbol{n}_{\ast}$.
@@
$=$ $\frac{(d-1)}{d}[1-L\gamma\omega],$
where in the the second equality we have used Eq. (2), and the third equality follows from the fact that the optimum detection happens when ${\textrm{Tr}}[\mathcal{O}^{(\alpha)}]$ takes its maximum value $d$, i.e., when $\mathcal{O}^{(\alpha)}={\mathbb{I}}_{d}$ for all $\alpha$. It turns out that $W$ can detect the entanglement of isotropic states when $\omega>\frac{1}{L\gamma}$. Accordingly, entanglement detection depends on the product of two parameters $L$ and $\gamma$: how many pairs of MEBs are used to detect entanglement, and how effective the measurements are. This provides a tradeoff relation between the number of measurements and their effectiveness in the sense that the more incompatible the measurements are, the fewer measurements are needed for detection. Clearly, complete entanglement detection is possible when a full set of $d+1$ measurements on MUBs are performed.
Figure 5 illustrates this quantity-quality tradeoff in terms of the index $\mu$ for $d=3$ and different values of $L=2,3,4$. In this case, $\lambda_{\max}(N)=\sqrt{(3\mu-1)/2}$, as such, the constraint (49) leads to $\mu<(2/(L-1)^{2}+1)/3$. This yields the bounds $\mu<1$, $\mu<1/2$, and $\mu<11/27$, for $L=2$, $L=3$, and $L=4$, respectively. In all cases, the solid curves represent $(L\gamma)^{-1}$ when MEBs are used to construct witnesses. The dashed lines, on the other hand, capture the bounds when a set of $L$ MUBs are used to construct witnesses, so that $\gamma=1$. Not surprisingly, for a fixed $L$ the MUBs-based witness is finer than the MEBs-based witness, however, if a MEBs-based witness employs a larger set of measurements, it can surpass the MUBs-based witness in detecting entanglement, particularly within the specific index $\mu$ of the underlying distribution $q$ of the MEBs.
Figure 5: Plots of $(L\gamma)^{-1}$ in terms of index $\mu$ for $d=3$ and different values of $L=2,3,4$. For equal $L$s, the MUBs-based witness (dashed lines) always captures a wider range of states relative to the MEBs-based witness (solid curves). However, if a MEBs-based witness employs a larger set of measurements, it can surpass the MUB-based witness in detecting entanglement, particularly within low values of the index $\mu$. As discussed in the main text, for $L=4$, the plot is restricted to $\mu\in[1/3,11/27)$, due to the imposed limit on $\mu$.
The inequality $\lambda_{\max}(N)<1/(L-1)$ imposes a tradeoff relation on the MEBs, in the sense that for a given set of MEBs associated with a probability distribution $q=(q_{0},\ldots,q_{d-1})$, only $L<1/\lambda_{\max}(N)+1$ of them are relevant for entanglement detection, even if more than $L$ could be constructed. The constraint does not make sense for MUBs, as $N_{k,d-k}=0$ for the uniform distribution $q=(1/d,\ldots,1/d)$, so they are already optimal. For MEBs, however, it arises from the possibility of linear dependence among the maximally entangled states in Eq. (57), which renders certain bases redundant for the entanglement detection. In view of this, the significance of MEBs is pronounced when their distribution is sufficiently close to the uniform distribution. Despite the constraint imposed on the MEBs, they may still be valuable in certain dimensions, such as $d=6$, where the maximum number of $3$ MUBs has been identified. This raises the question that, is it possible for a set of at least $4$ MEBs to exist in dimension $d=6$? and, if such a set is found, can it be sufficiently close to the uniform distribution such that the constraint does not exclude any of the bases? This possibility merits further investigation.
@@ V Conclusion
In this work, we introduced the concept of mutually equi-biased bases (MEBs), a generalization of the mutually unbiased bases (MUBs). For such bases, we relaxed the condition of mutual maximum unpredictability that holds for MUBs; instead, we associated a probability distribution $q=(q_{0},\ldots,q_{d-1})$ with each of the MEBs. This distribution captures the predictability of the outcomes of a measurement in one basis when the system is prepared in a state from the other basis. However, the information extracted from the measurement is unbiased, meaning that it cannot reveal which state from which basis the system is prepared in, and which basis is used for the measurement.
We derived a complete set of $d+1$ MEBs for $d=2$ and $d=3$, and found that although for $d=2$ a complete set of three MEBs exist for all values of the index $\mu=\sum_{k=0}^{d-1}q_{k}^{2}$, for $d\geq 3$ this is not the case. More precisely, for $d=3$, the constraints imposed by the mutual equi-biasedness limits the range of the index $\mu$ to the feasible region $1/3\leq\mu\leq 1/2$ within which a set of four MEBs exists. Furthermore, in addition to this restriction on $\mu$, for a given $\mu$ in the feasible region the components of the probability distribution $q=(q_{0},q_{1},q_{2})$ are also limited to some specific discrete points; only one probability distribution qq—up to permutation of its components—can generate a complete set of four MEBs.
For a qudit system, we derived an inequality for the probabilities of projective measurements in MEBs. We further used the inequality to derive an entropic uncertainty inequality in MEBs. As expected, the maximal incompatibility inherent to MUBs breaks down in the case of MEBs of the same size. The inequality for probabilities is then used to obtain a class of positive maps, from which we constructed entanglement witnesses. Using the MEBs for entanglement detection faces a constraint; depending on the underlying probability distribution $q$, the number of bases that can be employed from a set of $L$ MEBs is subject to a limit. This reveals a key distinction between MEBs and MUBs in such applications.
@@
These results provide a deeper understanding of the structure and properties of MEBs, offering new tools for entanglement detection and quantum measurement theory. The challenging dimension $d=6$ is currently under investigation to determine whether there exist values of the index $\mu$ for which one can construct a set of at least four MEBs—exceeding a set of three MUBs currently known for this dimension. We hope that this work will serve as a stepping stone for further research, ultimately advancing our understanding of MUBs. In particular, for dimensions where no more than three MUBs are known, even a set of four MEBs with $\mu>1/d$ could shed light on the challenging open problem of constructing complete set of MUBs.